mr. santos is 25years older than his son. how old is each, if the father is five years more than three times as old as his son?
Answer by amoresroy(361)

mr. santos is 25years older than his son. how old is each, if the father is five years more than three times as old as his son?
Let x = age of son
x+25 = age of father
x+25 = 3x + 5
2x = 20
x = 10
x+25 = 35
The son is 10 years old and the father is 35 years old.